OpenStudy (anonymous):

two forces of magnitude 6N and 4N act on a body.What is the ratio of the acceleration produced when these forces are acting in the same direction to the acceleration when oppositely directed?

OpenStudy (anonymous):

(6+4)/(6-4) = 10/2 = 5/1

OpenStudy (anonymous):

In same direction, it adds, in opposite it substracts. Hence, asnwer is 5

OpenStudy (anonymous):

thank you!is there any formula for this problem?